OpenStudy (anonymous):

Change –5x + 2y = 12 into slope-intercept form. I can get that it is Y = ?x + 6.

OpenStudy (callisto):

Slope-intercept is y = mx + c, where m is the slope and c is the y-intercept. Basically, all you have to do is the make y as the subject for the equation. Do you know how to make y as the subject?

OpenStudy (anonymous):

Is the M x a fraction?

OpenStudy (callisto):

Yup, for your question, it is.

OpenStudy (anonymous):

is it negative?

OpenStudy (callisto):

Not really... What have you got?

OpenStudy (anonymous):

Y = 5/2x +6

OpenStudy (callisto):

Bingo! :) If you mean \(y=\frac{5}{2}x+6\)

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Then, you got it :)